OpenStudy (anonymous):

Factor the trinomial S^2-10s+16

OpenStudy (anonymous):

s^2 - 10s +16 =s^2-8s-2s+16 =s(s-8)-2(s-8) =(s-8)(s-2)

OpenStudy (anonymous):

hmm .. ok so we are on the same topic..... when you factorize the answer wouldd be (?,? (?,?)

OpenStudy (anonymous):

the best way is to look at each variable...... s^2 is s x s that's the first number.... the